# Functional-Group-Based Diffusion for Pocket-Specific Molecule Generation and Elaboration

### v1

@@ 1 Introduction
Recently, a number of methods of AI-assisted drug design (AIDD) on pocket-specific molecule generation are emerging (peng2022pocket2mol, ; liu2022graphbp, ; tan2022target, ; luo20213d, ; guan3d, ; schneuing2022structurebased, ; diffbp, ), thanks to developments in deep generative models doersch2016tutorial ; goodfellow2014generative ; dinh2016density ; kingma2018glow ; nielsen2020survae ; bengio2021gflownet and graph neural networks (GNNs) gcn ; mpnn ; liu2021dig ; wu2021selfsuper ; liu2021spherical . These methods usually focus on atom-level generation, which first generates atom types and positions and then assembles atoms into molecules that can bind to the protein pockets. Although significant progress has been made, they are still weak in two aspects. For one thing, it is hard for them to generate realistic functional groups that contribute pharmacological effects to the target as classical CADD methods are able to. It is shown that generating benzene rings is uncommon compared with the reference molecules, not to mention some large functional groups with complex structure constraints such as purine, indole, etc (Table. 1 in Sec. 5.2). For another, trade-offs between efficiency and sufficiency of protein context place them in a dilemma. For example, TargetDiff guan3d employs sufficient protein context, which disassembles the amino acids into atoms, but leads to inefficiency due to a large node number (412.14 on average) in GNN’s message passing. DiffSBDD schneuing2022structurebased simplifies the representation of protein context, by only using ${\mathrm{C}_{\alpha}}$’s positions and residue types, resulting in a reduction of GNNs’ node number (68.10 on average) but the insufficiency of context information.
@@ 2 Problem Statement
For a binding system composed of a protein-molecule pair (also called protein-ligand pair) as $\mathcal{C}$, which contains $N_{\mathrm{aa}}$ amino acids of proteins and $N_{\mathrm{fg}}$ functional groups and $N_{\mathrm{at}}$ single atoms of molecules, we represent the index set of the molecule’s single atoms as $\mathcal{I}_{\mathrm{at}}$, functional groups as $\mathcal{I}_{\mathrm{fg}}$, and the protein’s amino acids as $\mathcal{I}_{\mathrm{aa}}$, where $|\mathcal{I}_{\mathrm{fg}}|=N_{\mathrm{fg}}$, $|\mathcal{I}_{\mathrm{at}}|=N_{\mathrm{at}}$ and $|\mathcal{I}_{\mathrm{aa}}|=N_{\mathrm{aa}}$. Note that a molecule can be disassembled into functional groups and single atoms other than functional groups, which we also call linkers. For a protein, the amino acids can be represented by its type, $\mathrm{C}_{\alpha}$ atom coordinate, and the orientation, denoted as $s_{i}\in\{1,\ldots,20\}$, $\bm{x}_{i}\in\mathbb{R}^{3}$, $\bm{O}_{i}\in\mathrm{SO(3)}$, where $i\in\mathcal{I}_{\mathrm{aa}}$. For a molecule, assuming there are $M_{\mathrm{fg}}$ and $M_{\mathrm{at}}$ possible types in total functional groups and linker atoms respectively, the functional groups can be represented as the three elements if the inter-relative positions are fixed, as $s_{j}$, $\bm{x}_{j}$ and $\bm{O}_{j}$, where $s_{j}\in\{21,\ldots,21+M_{\mathrm{fg}}\}$ is the type, $\bm{x}_{j}\in\mathbb{R}^{3}$ is the predefined center atom’s coordinate, and $\bm{O}_{j}\in\mathrm{SO(3)}$ can also be obtained in the same way as amino acids (See Sec. 5.1 and Appendix. C.) for $j\in\mathcal{I}_{\mathrm{fg}}$; And its linkers can be represented as $s_{k}$, $\bm{x}_{k}$ and $\bm{O}_{k}$, with $s_{k}\in\{22+M_{\mathrm{fg}},\ldots,22+M_{\mathrm{fg}}+M_{\mathrm{at}}\}$, $k\in\mathcal{I}_{\mathrm{at}}$ and $\bm{O}_{k}=\mathrm{diag}\{1,1,1\}=\bm{I}$.
@@ 3 Related Work
Success in 3D molecule generation and an increasing amount of available structural data of protein and molecules raises scientific interests in structure-based drug design (SBDD), which aims to generate both 2D molecule graphs and 3D structures conditioned on target protein structure as contextual constraints. For example, LiGAN masuda2020generating and 3DSBDD luo20213d are grid-based models which predict whether the grid points are occupied by specific atoms. By harnessing 3D equivariant graph neural networks, Pocket2Mol peng2022pocket2mol and GraphBP liu2022graphbp generate atoms auto-regressively and model the probability of the next atom’s type as discrete categorical attribute and position as continuous geometry. Recently, utilizing the diffusion denoising probability models ho2020denoising ; song2019generative ; song2020score ; emiel2021edm ; karras2022elucidating , a series of SBDD methods generate ligands conditioned on the target pockets at full atom levels guan3d ; schneuing2022structurebased ; diffbp .
@@ 4.1 Diffusion Models
Let $\bm{s}_{j}^{t}$ as the one-hot encoding of the type of a single functional group or linker. The forward process randomly transits $s_{t}$ into the absorbing state $K$ ($K=23+M_{\mathrm{fg}}+M_{\mathrm{at}}$) with
@@
By defining the center atom in a functional group as shown in Appendix. C, and itself as the center in a linker, the Cartesian coordinate of center nodes $\bm{x}_{j}$ represents its position. The forward transition distributions followed by DDPM ho2020denoising read
@@ 4.2 Parametrization of Denoisers with Neural Networks
In order to decrease the computational complexity, we denote the protein context at amino-acid levels. Besides the amino acid types, $\mathrm{C}_{\alpha}$ atom coordinate and the orientation, each atom’s coordinates in the local system and three torsion angles including angles around ‘$\mathrm{N}$-$\mathrm{C}_{\alpha}$’ bond, ‘$\mathrm{C}_{\alpha}$-$\mathrm{C}$’ bond and ‘$\mathrm{C}$-$\mathrm{N}$’ bond, are also used as intra-amino-acid features, which are concatenated and embedded by an MLP to create the intra-amino-acid embedding vector $\bm{e}_{i}$. For inter-amino-acid features, the pair of amino acid types, sequential relationships (if the two amino acids are adjacent in the protein sequence), pairwise distances between $\mathrm{C}_{\alpha}$ and inter-residue dihedrals are all embedded as inter-amino-acid embedding vector $\bm{z}_{i,j}$ with $i,j\in\mathcal{I}_{\mathrm{aa}}$. Note that these embedding vectors are all translational and rotational invariant (Appendix. B).
@@
The output of $G$ means predicting the coordinate deviations in the local coordinate systems and then projecting it into the global one; In $H$, $\mathrm{MLP}_{H}$ first predicts a vector in Lie group $\mathfrak{so}(3)$ and the exponential map on SO(3) converts it into a rotation matrix. The updating process ensures the equivariance and the invariance of the transition distributions in Proposition 1, which is proved in kofinas2022rototranslated and luo2022antigenspecific . However, since $\bm{O}_{i}^{t}=\bm{I}$ for any $i\in\mathcal{I}_{\mathrm{at}}$, the equivariance of ${G}$ on linkers will not be preserved, so we instead use EGNN victor2022egnn to get the output of $G$ (Appendix. B).
@@ 5 Experiments
In the experiments, we use CrossDocked2020crossdocked2020 for evaluation. In the prevailing works, they focus on generating molecules at the atom level, differing from our functional-group-based generation, so our first target is to divide the molecules into functional groups and linkers. We use EFGsErtl2017AnAT to segment molecules. We select the 25 functional groups that appear most frequently with stable structures, which are partly shown in Figure. 2. For some functional groups, chirality exists in their structures, and we treat them as two functional groups. Therefore, we finally build up a dataset as a corpus including 25 functional groups (two of which have chirality) for Crossdocked2020, with their intra-structures fixed as rigid bodies. Details are given in Appendix. C. For linkers, we choose $\{\mathrm{B},\mathrm{C},\mathrm{N},\mathrm{O},\mathrm{F},\mathrm{P},\mathrm{S},\mathrm{Cl},\mathrm{Br},\mathrm{I}\}$ as representative heavy atoms. After the processing, we obtain $M_{\mathrm{fg}}=25$ and $M_{\mathrm{at}}=10$ in our experiments. Besides, by the fragment-linker designation of a binding graph, the node number is reduced to 53.62 on average in GNN’s message-passing.
@@ 5.2 Pocket-Specific Molecule Generation
Table 1: ‘Ratio’ of the top ten functional groups with the highest frequency in Crossdocked2020. ‘Ref.’ is calculated in the training set. MAE is obtained between rows of ‘Ref.’ and different methods’ ‘Ratio’ and JSD is calculated by ‘Freq’, which is in detail given in Appendix. D ‘Ratio’ in bold means it is the closest to ‘Ref.’, and MAE/JSD in bold means it is the lowest.
@@
We first analyze the functional groups of generated molecules by different methods. We show the ‘Ratio’ and ‘Freq’ of the included functional groups partly, where ‘Ratio’ means how many specified functional groups are in each molecule on average, and ‘Freq’ means the statistical frequency of occurrence of the specified functional group in the generated molecules. Mean absolute error (MAE) as Overall metrics is calculated according to reference ‘Ratio’ and generated ‘Ratio’; Jensen-Shannon divergence (JSD) is calculated according to reference ‘Freq’ and generated ‘Freq’. The smaller MAE and JSD are, the better performance the method achieves. Table. 1 shows the metrics of different methods’ generated molecules, demonstrating D3FG’s superiority in generating molecules with realistic drug structures since distributions of the functional groups in generated molecules are more similar to the real drug molecules’. Secondly, we analyze the bond distance distribution and atom type distribution. JSDs are calculated between the distributions of bond distance for reference vs. generated molecules. ‘-’, ‘=’,‘:’ represents single, double, and aromatic bonds, respectively. Besides, we report MAE on reference ‘Ratio’ vs. generated ‘Ratio’ and JSD on reference ‘Freq’ vs. generated ‘Freq’ based on atom type distribution. Table. 2, 3 and Figure. 3 gives further details. We found that D3FG(Stage) outperforms D3FG(Joint) in generating more realistic molecules, and has competitive performance with TargetDiff.
@@ Introduction.
We here consider two schemes for molecule elaboration, the first is to remove the functional groups with the highest scores, and elaborate the remaining fragments by replacing them with the functional group generated by D3FG. We write this as D3FG(EHot). The second, on the other hand, replaces the functional groups of the lowest scores, which is represented by D3FG(ECold). We report the D3FG’s elaborated molecules by the two schemes in Table. 4. It shows that D3FG(EHot) tends to generate more molecules with higher affinity, while tiny differences in binding affinity between molecules elaborated by D3FG(ECold) and the raw references. Besides, on other chemical properties, the elaborated molecules are very close to the raw references, since the differences lie only in one single functional group, and the major molecular skeletons are almost identical.
@@ Acknowledgements
SE(3) group as a roto-translation group in $\mathbb{R}^{3}$, can be divided into two groups: SO(3) as the rotation group and T(3) as the translation group. For $\bm{x}\in\mathbb{R}^{3}$, and $g=r+v$ with $g\in\mathrm{SE(3)},r\in\mathrm{SO(3)},v\in\mathrm{T(3)}$, $T_{g}(\bm{x})=T_{r+v}(\bm{x})=T_{v}\circ T_{r}(\bm{x})$
@@
Appendix B Method Details

### v3

@@ 1 Introduction
Artificial intelligence has already achieved outstanding success in protein design (alphadesign, ; gao2022alphadesign, ; huang2023dataefficient, ; wu2022survey, ; gao2022pifold, ; tan2022generative, ) and thus led to a new round of attention in drug design focused on AI-assisted drug design (AIDD). A number of methods of AI-assisted drug design on pocket-specific molecule generation are emerging (peng2022pocket2mol, ; liu2022graphbp, ; tan2022target, ; luo20213d, ; guan3d, ; schneuing2022structurebased, ; diffbp, ), thanks to developments in deep generative models doersch2016tutorial ; goodfellow2014generative ; dinh2016density ; kingma2018glow ; nielsen2020survae ; bengio2021gflownet and graph neural networks (GNNs) gcn ; mpnn ; liu2021dig ; wu2021selfsuper ; liu2021spherical ; wu2021graphmixup ; wu2023beyond ; wu2022kdist ; wu2023extracting ; wu2023quantifying . These methods usually focus on atom-level generation, which first generates atom types and positions and then assembles atoms into molecules that can bind to the protein pockets. Although significant progress has been made, they are still weak in two aspects. For one thing, it is hard for them to generate realistic functional groups that contribute pharmacological effects to the target as classical CADD methods are able to. It is shown that generating benzene rings is uncommon compared with the reference molecules, not to mention some large functional groups with complex structure constraints such as purine, indole, etc (Table. 1 in Sec. 5.2). For another, trade-offs between efficiency and sufficiency of protein context place them in a dilemma. For example, TargetDiff guan3d employs sufficient protein context, which disassembles the amino acids into atoms, but leads to inefficiency due to a large node number (412.14 on average) in GNN’s message passing. DiffSBDD schneuing2022structurebased simplifies the representation of protein context, by only using ${\mathrm{C}_{\alpha}}$’s positions and residue types, resulting in a reduction of GNNs’ node number (68.10 on average) but the insufficiency of context information.
@@ 2 Problem Statement
For a binding system composed of a protein-molecule pair (also called protein-ligand pair) as $\mathcal{C}$, which contains $N_{\mathrm{aa}}$ amino acids of proteins and $N_{\mathrm{fg}}$ functional groups and $N_{\mathrm{at}}$ single atoms of molecules, we represent the index set of the molecule’s single atoms as $\mathcal{I}_{\mathrm{at}}$, functional groups as $\mathcal{I}_{\mathrm{fg}}$, and the protein’s amino acids as $\mathcal{I}_{\mathrm{aa}}$, where $|\mathcal{I}_{\mathrm{fg}}|=N_{\mathrm{fg}}$, $|\mathcal{I}_{\mathrm{at}}|=N_{\mathrm{at}}$ and $|\mathcal{I}_{\mathrm{aa}}|=N_{\mathrm{aa}}$. Note that a molecule can be disassembled into functional groups and single atoms other than functional groups, which we also call linkers. For a protein, the amino acids can be represented by its type, $\mathrm{C}_{\alpha}$ atom coordinate, and the orientation, denoted as $s_{i}\in\{1,\ldots,20\}$, $\bm{x}_{i}\in\mathbb{R}^{3}$, $\bm{O}_{i}\in\mathrm{SO(3)}$, where $i\in\mathcal{I}_{\mathrm{aa}}$. For a molecule, assuming there are $M_{\mathrm{fg}}$ and $M_{\mathrm{at}}$ possible types in total functional groups and linker atoms respectively, the functional groups can be represented as the three elements if the inter-relative positions are fixed, as $s_{j}$, $\bm{x}_{j}$ and $\bm{O}_{j}$, where $s_{j}\in\{21,\ldots,21+M_{\mathrm{fg}}\}$ is the type, $\bm{x}_{j}\in\mathbb{R}^{3}$ is the predefined center atom’s coordinate, and $\bm{O}_{j}\in\mathrm{SO(3)}$ can also be obtained in the same way as amino acids (See Sec. 5.1 and Appendix. D.) for $j\in\mathcal{I}_{\mathrm{fg}}$; And its linkers can be represented as $s_{k}$, $\bm{x}_{k}$ and $\bm{O}_{k}$, with $s_{k}\in\{22+M_{\mathrm{fg}},\ldots,22+M_{\mathrm{fg}}+M_{\mathrm{at}}\}$, $k\in\mathcal{I}_{\mathrm{at}}$ and $\bm{O}_{k}=\mathrm{diag}\{1,1,1\}=\bm{I}$.
@@ 3 Related Work
Fragment-based drug design.
Previously, works on fragment-based molecule generation are proposed. For example, JT-VAE jtvae generates a tree-structured scaffold over chemical substructures and combines them into a 2D-molecule. PS-VAE psvae can automatically discover frequent principal subgraphs from the dataset, and assemble generated subgraphs as the final output molecule in 2D. Further, DeepFrag deepfrag predicts fragments conditioned on parents and the pockets, SQUID squid generates molecules in a fragment level conditioned on molecule’s shapes. FLAGflag auto-regressively generates fragments as motifs based on the protein structures in 3D.
@@
Success in 3D molecule generation and an increasing amount of available structural data of protein and molecules raises scientific interests in structure-based drug design (SBDD), which aims to generate both 2D molecule graphs and 3D structures conditioned on target protein structure as contextual constraints. For example, LiGAN masuda2020generating and 3DSBDD luo20213d are grid-based models which predict whether the grid points are occupied by specific atoms. By harnessing 3D equivariant graph neural networks, Pocket2Mol peng2022pocket2mol and GraphBP liu2022graphbp generate atoms auto-regressively and model the probability of the next atom’s type as discrete categorical attribute and position as continuous geometry. FLAGflag is also fragment-based, but still generates motifs in an auto-regressive way. Recently, utilizing the diffusion denoising probability models ho2020denoising ; song2019generative ; song2020score ; emiel2021edm ; karras2022elucidating , a series of SBDD methods generate ligands conditioned on the target pockets at full atom levels guan3d ; schneuing2022structurebased ; diffbp .
@@ 4 Method
D3FG firstly decomposes molecules into two categories of components: functional groups and linkers, and them use the diffsion generative model to learn the type and geometry distributions of the components. In this section, we describe the D3FG by four parts: (i) the diffusion model as the generative framework, in which the three variables are generated; (ii) the denoiser parameterized by graph neural networks, satisfying certain symmetries so that the generative model is SE-(3) invariant; (iii) the sampling process in which the molecules are generated by the trained models; (iv) the further problems resulted from heterogenous graph with two solutions.
@@ 4.1 Diffusion Models
Let $\bm{s}_{j}^{t}$ as the one-hot encoding of the type of a single functional group or linker. The forward process followed by D3PM austin2023structured randomly transits $s^{t}_{j}$ into the absorbing state $K$ ($K=23+M_{\mathrm{fg}}+M_{\mathrm{at}}$) with
@@
By defining the center atom in a functional group as shown in Appendix. D, and itself as the center in a linker, the Cartesian coordinate of center nodes $\bm{x}_{j}$ represents its position. The forward transition distributions followed by DDPM ho2020denoising read
@@ 4.2 Parametrization of Denoisers with Neural Networks
In order to decrease the computational complexity, we denote the protein context at amino-acid levels. Besides the amino acid types, $\mathrm{C}_{\alpha}$ atom coordinate and the orientation, each atom’s coordinates in the local system and three torsion angles including angles around ‘$\mathrm{N}$-$\mathrm{C}_{\alpha}$’ bond, ‘$\mathrm{C}_{\alpha}$-$\mathrm{C}$’ bond and ‘$\mathrm{C}$-$\mathrm{N}$’ bond, are also used as intra-amino-acid features, which are concatenated and embedded by an MLP to create the intra-amino-acid embedding vector $\bm{e}_{i}$. For inter-amino-acid features, the pair of amino acid types, sequential relationships (if the two amino acids are adjacent in the protein sequence), pairwise distances between $\mathrm{C}_{\alpha}$ and inter-residue dihedrals are all embedded as inter-amino-acid embedding vector $\bm{z}_{i,j}$ with $i,j\in\mathcal{I}_{\mathrm{aa}}$. Note that these embedding vectors are all translational and rotational invariant (Appendix. C).
@@
The output of $G$ means predicting the coordinate deviations in the local coordinate systems and then projecting it into the global one; In $H$, $\mathrm{MLP}_{H}$ first predicts a vector in Lie group $\mathfrak{so}(3)$ and the exponential map on SO(3) converts it into a rotation matrix. The updating process ensures the equivariance and the invariance of the transition distributions in Proposition 1, which is proved in kofinas2022rototranslated and luo2022antigenspecific . However, since $\bm{O}_{i}^{t}=\bm{I}$ for any $i\in\mathcal{I}_{\mathrm{at}}$, the equivariance of ${G}$ on linkers will not be preserved, so we instead use EGNN victor2022egnn to get the output of $G$ (Appendix. C).
@@ 5 Experiments
In the experiments, we use CrossDocked2020crossdocked2020 for evaluation. In the prevailing works, they focus on generating molecules at the atom level, differing from our functional-group-based generation, so our first target is to divide the molecules into functional groups and linkers. We use EFGsErtl2017AnAT to segment molecules. We select the 25 functional groups that appear most frequently with stable structures, which are partly shown in Figure. 2 For some functional groups, chirality exists in their structures, and we treat them as two functional groups. As a result, we finally build up a dataset as a corpus including 27 functional groups (two of the 25 have chirality) for Crossdocked2020, with their intra-structures fixed as rigid bodies, and assure that most molecules can be decomposed into the substructures in our corpus datasets. Details are given in Appendix. D. For linkers, we choose $\{\mathrm{B},\mathrm{C},\mathrm{N},\mathrm{O},\mathrm{F},\mathrm{P},\mathrm{S},\mathrm{Cl},\mathrm{Br},\mathrm{I}\}$ as representative heavy atoms. After the processing, we obtain $M_{\mathrm{fg}}=27$ and $M_{\mathrm{at}}=10$ in our experiments. Besides, by the fragment-linker designation of a binding graph, the node number is reduced to 53.62 on average in GNN’s message-passing.
@@ 5.2 Pocket-Specific Molecule Generation
Table 1: ‘Ratio’ of the top ten functional groups with the highest frequency in Crossdocked2020. ‘Ref.’ is calculated in the training set. MAE is obtained between rows of ‘Ref.’ and different methods’ ‘Ratio’ and JSD is calculated by ‘Freq’, which is in detail given in Appendix. E ‘Ratio’ in bold means it is the closest to ‘Ref.’, and MAE/JSD in bold means it is the lowest.
@@
We first analyze the functional groups of generated molecules by different methods. We show the ‘Ratio’ and ‘Freq’ of the included functional groups partly, where ‘Ratio’ means how many specified functional groups are in each molecule on average, and ‘Freq’ means the statistical frequency of occurrence of the specified functional group in the generated molecules. Mean absolute error (MAE) as overall metrics is calculated according to reference ‘Ratio’ and generated ‘Ratio’; Jensen-Shannon divergence (JSD) is calculated according to reference ‘Freq’ and generated ‘Freq’. The smaller MAE and JSD are, the better performance the method achieves. Table. 1 shows the metrics of different methods’ generated molecules, demonstrating D3FG’s superiority in generating molecules with realistic drug structures since distributions of the functional groups in generated molecules are more similar to the real drug molecules’. Secondly, we analyze the atom type, bond distance, bond angle, dihedral and atom type distribution. JSDs are calculated between the distributions of bond distance for reference vs. generated molecules. ‘-’, ‘=’,‘:’ represents single, double, and aromatic bonds, respectively. Besides, we report MAE on reference ‘Ratio’ vs. generated ‘Ratio’ and JSD on reference ‘Freq’ vs. generated ‘Freq’ based on atom type distribution. Table. 2, 3 and Figure. 3 gives details atom type analysis. For other geometries, please refer to Appendix. E. We found that D3FG(Stage) outperforms D3FG(Joint) in generating more realistic molecules, and has competitive performance with TargetDiff.
@@ Introduction.
We here consider two schemes for molecule elaboration, the first is to remove the functional groups with the highest scores, and elaborate the remaining fragments by replacing them with the functional group generated by D3FG. We write this as D3FG(EHot). The second D3FG(ECold), on the other hand, replaces the functional groups of the lowest scores). We report the D3FG’s elaborated molecules by the two schemes in Table. 4. It shows that D3FG(EHot) tends to generate more molecules with higher affinity, while molecules elaborated by D3FG(ECold) show tiny differences in binding affinity from the raw references. Besides, on other chemical properties, the elaborated molecules are very close to the raw references, since the differences lie only in one single functional group, and the major molecular skeletons are almost identical.
@@ Acknowledgements
SE(3) group as a roto-translation group in $\mathbb{R}^{3}$, can be divided into two groups: SO(3) as the rotation group and T(3) as the translation group. For $\bm{x}\in\mathbb{R}^{3}$, and $g=r+v$ with $g\in\mathrm{SE(3)},r\in\mathrm{SO(3)},v\in\mathrm{T(3)}$, $T_{g}(\bm{x})=T_{r+v}(\bm{x})=T_{v}\circ T_{r}(\bm{x})$. The proof is inspired by GeoDiff and EF [50, 85].
@@
Appendix B Model Comparisons.
With Pocket2Mol and GraphBP.
Pocket2Mol and GraphBP are all auto-regressive models, which violate the physical rules from the perspective of energy[15], while the diffusion models which consider the global interactions are a solution to the problem. Besides, to decide which atom will be added a bond with the next atom, Pocket2Mol needs a classifier to predict the focal atom, so that the training and the prediction are not end-to-end. Finally, D3FG is a functional-group- or fragment-based method, while these two models generate molecules at the atom level.
With DiffSBDD and TargetDiff.
These two methods are diffusion-based, which considers the global interactions between atoms. In the atom type generation, DiffSBDD uses continuous latent spaces generation, while Targetdiff diffuses the atom types in a uniform distribution. These two models all generate molecules at the atom level, while D3FG generates at the fragment level, and besides the types and positions, the orientations of the functional groups are generated.
With FLAG.
It is still an auto-regressive model and does not support end-to-end generation since the focal atom needs to be predicted. Besides, we defined our functional groups as rigid bodies with stable intra-structures, while the motifs in FLAG are 2D smiles, with the structures generated by RDKIT. The 3D structures of functional groups are obtained by the training set in D3FG, thus avoiding the problem of distribution shift of FLAG since the training/test motif substructures may not match the RDKIT’s generation.
With DeepFrag.
It is a model for fragment-based lead optimization, in which the protein and the parent is used as condition and the fragment type is predicted by the model. In this way, it is more like an elaboration task defined in D3FG. Here we point out several advantages of D3FG(EHot/Cold) over Deepfrag. First, Deepfrag just predicts the fragment type, without 3D structures, so that the problem of equivariance is not included, but D3FG generates 3D positions of the molecules. Second, although the type is the SE(3)-invariant variable, the model uses CNNs as the model backbone rather than EGNN, so the invariance can also not be preserved. Instead, D3FG assures physical symmetries by using EGNN. Finally, D3FG is a generative model with stochasticity, while Deepfrag only gives the probability of the fragment types, as a discriminative model.
Appendix C Method Details